# Position Paper: Against Spurious Sparks—Dovelating Inflated AI Claims [Uncaptioned image]

### v2

@@ 3.1 Spurious Relationships
A major contributor is that humans struggle to tell the difference between random and non-random sequences (Falk and Konold, 1997), and to generate sequences that appear random (Ladouceur et al., 1996). A common issue is a lack of expectation that randomness that hints towards a causal relationship, such as correlations, will still appear at random. This leads even those trained in statistics and probability to perceive illusory correlations, correlations of inflated magnitude (see Nickerson (1998)), or causal relationships in data that is randomly generated (Zgraggen et al., 2018).
@@ 3.2 Anthropomorphism
Research on anthropomorphism has repeatedly shown the human tendency to attribute human-like characteristics to non-human agents and/or objects. These might include the weather and other natural forces, pets and other animals, gadgets and other pieces of technology (Epley et al., 2007). Formally studied as early as 1944, (Heider and Simmel, 1944) observed that humans can correctly interpret a narrative whose characters are abstract 2D shapes, but also that humans interpreted random movements of these shapes as having a human-like narrative. Relevant to AI and the degree to which it resembles AGI, anthropomorphizing may occur independently of whether such judgments are accurate, and as a matter of degree: at the weaker end, one may employ anthropomorphism as a metaphorical way of thinking or explaining, and at the stronger end one may attribute human emotions, cognition, and intelligence to AI systems. As Epley et al. (2007) note, literature has shown that even weak metaphorical anthropomorphism may affect how humans behave towards non-human agents.
@@
Modern anthropomorphism theory suggests there are three key components, one of which is a cognitive feature, and two of which are motivations. The first involves the easy availability of our experiences as heuristics that can be used to explain external phenomena: “…knowledge about humans in general, or self-knowledge more specifically, functions as the known and often readily accessible base for induction about the properties of unknown agents" (p.866) (Epley et al., 2007). Thus, our experience as humans is an always-readily-available template to interpret the world, including non-human agent behaviors.
This may be more so when the behaviors of that agent are made to resemble humans, which can be a benefit to the second key component of the theory: a motivational state to anthropomorphize among individuals experiencing loneliness, social isolation, or otherwise seeking social connection (Epley et al., 2007; Waytz et al., 2010).
@@
The third component is the motivation as a human to be competent (effectance motivation). This is most relevant to this discussion, as it describes the need to effectively interact with our environments, including the technologies of the day (Epley et al., 2007). When confronted with an opaque technology, a person may interpret its behaviors using the most readily available template at hand, namely their personal human experience, in order to facilitate learning (Epley et al., 2007; Waytz et al., 2010). Perceiving human characteristics, motivations, emotions, and cognitive processes from one’s own experiences in a technology e.g. an AI chatbot, allows for a ready template of comparison at the very least, and possibly an increase in ability to make sense of, and even predict, the agent’s behaviors. This may include being placed in a position to master a certain technology, whether by incentives to learn, or fear of poor outcomes should one not manage to learn.
These pressures extend to AI experts, as well as laypersons. In both scholarly and commercial fields, AI experts face considerable pressure to demonstrate competence in their work. Citation metrics and scholarly publications remain the primary metric for tenure and promotion, e.g. (Alperin et al., 2019), and the number of publications in the AI field has boomed as evidenced by overall (preprint and peer-reviewed) scholarly publications [^5] https://ourworldindata.org/grapher/annual-scholarly-publications-on-artificial-intelligence?time=2010..2021  (Maslej et al., 2023).
@@
The adoption of techniques underlying technologies with the AI label, i.e. machine learning, has spread to fields beyond Computer Science, e.g. Astronomy, Physics, Medicine and Psychology [^6] Retrieved 23/01/23 using the search string ”TITLE-ABS-KEY ( ( machine AND learning ) OR ( artificial AND intelligence ) OR ai ) AND PUBYEAR $>$ 2009 AND PUBYEAR $<$ 2024 ” from the SCOPUS database. Outside of academia, the number of jobs requiring AI expertise increases rapidly, with demand for ‘Machine Learning’ skills clusters having increased over 500% from 2010 to 2020 (Maslej et al., 2023). Thus, according to theory, the pressure to demonstrate AI-competence is fertile ground for anthropomorphism to occur.
@@ 3.3 Confirmation Bias
Hypotheses in present-day AI research are often implicit. Generally, these hypotheses are framed simply as a system being more accurate or efficient, compared to other systems. Where other fields, such as medicine or quantitative social sciences, would further articulate expectations in e.g. assigning specific conditions and considering effect sizes assigned to each competing hypothesis, in computer science and AI this is typically not done. This also may have to do with much of the published work being more of an engineering achievement, rather than a true hypothesis test seeking to explain and understand the world. However, in discussions on emerging qualities like AGI, this engineering positioning gets muddier, and more formal hypothesis testing would be justifiable: either one interprets outputs as in support of hints towards AGI (the competing hypothesis), or as merely the result of an algorithm integrating qualities from the data it was trained on (the null hypothesis).
@@ 4 Conclusion
As discussed above, AI research and development outcomes can very easily be over-interpreted, both from a data perspective, and because of human biases and interests. Even academic researchers are not free from such biases. Thus, we call for the community to create explicit room for organized skepticism.
@@
For research that seeks to explain a phenomenon, clear hypothesis articulation and strong null hypothesis formulation will be needed. If claims of human-like or superhuman intelligence are made, these should be subject to severe tests (Claesen et al., 2022) that go beyond the display of surprise. Apart from focusing on getting novel improvements upon state-of-the-art published, organizing red-teaming activities as a community may help in incentivizing and normalizing constructive adversarial questioning. As the quest for AGI is so deeply rooted in human-like recognition, adding our voice to emerging calls to be vigilant in communication (Shanahan, 2024), we put in an explicit word of warning about the use of terminology. Many terms used in current A(G)I research (e.g. emergence, intelligence, learning, ‘better than human’ performance) have a common understanding in specialized research communities, but have bigger, anthropomorphic connotations in laypersons. In fictional media, depictions of highly intelligent AI have for long been going around. In a unique study of films featuring fictional robots, defined as "…an artificial entity that can sense and act as a result of (real-world or fictional) technology…", in the 134 most highly rated science-fiction movies on IMDB, 74 out of the 108 AI-robots studied had a humanoid shape, and 68 out of those had sufficient intelligence to interact at an almost human-level (Saffari et al., 2021). The authors identify human-like communication and the ability to learn as essential abilities in the depiction of AI agents in movies. They further show a common plot: humans perceive the AI agents as inferior, despite their possession of self-awareness and the desire to survive, which fuels the central conflict of the film, wherein humanity is threatened by AI superior in both intellect and physical abilities. Moreover, it is often noted that experts and fictional content creators interact, informing and inspiring each other (Saffari et al., 2021; Neri and Cozman, 2020).
@@
This image also permeates present-day non-fictional writings on AI, which often heavily use anthropomorphized language (e.g. “ever more powerful digital minds” in the ‘Pause Giant AI Experiments’ open letter (Future of Life Institute, 2023)). News outlets love a story that sells and so we witness examples of humans falling in love with their AI chatbots (Morrone, 2023; Steinberg, 2023). The same news outlets discuss the human-like responses of Microsoft’s Bing search engine, which had at that point recently been adopting GPT-4777A large multimodal language model from OpenAI https://openai.com/research/gpt-4. The article (Cost, 2023), states “As if Bing wasn’t becoming human enough” and goes on to claim it told them it loves them. Here, AI experts and influencers also have considerable influence on how the narrative unfolds on social media: according to Neri and Cozman (2020), actual AI-related harms did not trigger viral amplification (e.g. the death of an individual dying while a Tesla car was in autopilot, or the financial bankruptcy of a firm using AI technology to execute stock trades). Rather, potential risks expressed by someone perceived as having expertise and authority were amplified, such as statements made by Stephen Hawking during an interview in 2014.
@@ A.2.1 Data
To estimate economic growth, we will rely on a quarterly series of the real gross domestic product (GDP) provided by the Federal Reserve Bank of St. Louis. The data arrives in terms of levels of real GDP. In order to estimate growth, we transform the data using log differences. Since our yield curve data is daily, we aggregate it to the quarterly frequency by taking averages of daily yields for each maturity. We also standardize yields since deep learning models tend to perform better with standardized data [Michal S Gal, 2019]. Since COVID-19 was a substantial structural break in the time series, we also filter out all observations after 2018.
@@ A.2.2 Model
The encoder consists of a single fully connected hidden layer with 32 neurons and a hyperbolic tangent activation function. The bottleneck layer connecting the encoder to the decoder, is a fully connected layer with 6 neurons. The decoder consists of two fully connected layers, each with a hyperbolic tangent activation function: the first layer consists of 32 neurons and the second layer will have the same dimension as the input data. The output layer consists of a single neuron for our output variable, $g_{t}$. We train the model over 1,000 epochs to minimize mean squared error loss using the Adam optimizer [Kingma and Ba, 2017].
@@ A.2.3 Linear Probe
While the loss function applies most direct pressure on layers near the final output layer, any information useful for the downstream task first needs to pass through the bottleneck layer [Alain and Bengio, 2016]. On a per-neuron basis, the pressure to distill useful representation is therefore likely maximized there. Consequently, the bottleneck layer activations seem like a natural place to start looking for compact, meaningful representations of distilled information. We compute and extract these activations $A_{t}$ for all time periods $t=1,...,T$. Next, we use a linear probe to regress the observed yield curve factors on the latent embeddings. Let $Y_{t}$ denote the vector containing the two factors of interest in time $t$: $y_{t,l}$ and $y_{t,s}$ for the level and spread, respectively. Formally, we are interested in the following regression model: $p_{w}(Y_{t}|A_{t})$ where $w$ denotes the regression parameters. We use Ridge regression with $\lambda$ set to $0.1$. Using the estimated regression parameters $\hat{w}$, we then predict the yield curve factors: $\hat{Y}_{t}=\hat{w}^{\prime}A_{t}$.

### v3

@@ 4.1 Spurious Relationships
A major contributor is that humans struggle to tell the difference between random and non-random sequences (Falk & Konold, 1997), and to generate sequences that appear random (Ladouceur et al., 1996). A common issue is a lack of expectation that randomness that hints towards a causal relationship, such as correlations, will still appear at random. This leads even those trained in statistics and probability to perceive illusory correlations, correlations of inflated magnitude (see Nickerson (1998)), or causal relationships in data that is randomly generated (Zgraggen et al., 2018).
@@ 4.2 Anthropomorphism
Research on anthropomorphism has repeatedly shown the human tendency to attribute human-like characteristics to non-human agents and/or objects. These might include the weather and other natural forces, pets and other animals, gadgets and other pieces of technology (Epley et al., 2007). Formally studied as early as 1944, Heider & Simmel (1944) observed that humans can correctly interpret a narrative whose characters are abstract 2D shapes, but also that humans interpreted random movements of these shapes as having a human-like narrative. Relevant to AI and the degree to which it resembles AGI, anthropomorphizing may occur independently of whether such judgments are accurate, and as a matter of degree: at the weaker end, one may employ anthropomorphism as a metaphorical way of thinking or explaining, and at the stronger end one may attribute human emotions, cognition, and intelligence to AI systems. As Epley et al. (2007) note, literature has shown that even weak metaphorical anthropomorphism may affect how humans behave towards non-human agents.
@@
Modern anthropomorphism theory suggests there are three key components, one of which is a cognitive feature, and two of which are motivations. The first involves the easy availability of our experiences as heuristics that can be used to explain external phenomena: “…knowledge about humans in general, or self-knowledge more specifically, functions as the known and often readily accessible base for induction about the properties of unknown agents” (p.866 in Epley et al. (2007)). Thus, our experience as humans is an always-readily-available template to interpret the world, including non-human agent behaviors. This may be more so when the behaviors of that agent are made to resemble humans, which can be a benefit to the second key component of the theory: a motivational state to anthropomorphize among individuals experiencing loneliness, social isolation, or otherwise seeking social connection (Epley et al., 2007; Waytz et al., 2010).
@@
The third component is the motivation as a human to be competent (effectance motivation). This is most relevant to this discussion, as it describes the need to effectively interact with our environments, including the technologies of the day (Epley et al., 2007). When confronted with an opaque technology, a person may interpret its behaviors using the most readily available template at hand, namely their personal human experience, in order to facilitate learning (Epley et al., 2007; Waytz et al., 2010). Perceiving human characteristics, motivations, emotions, and cognitive processes from one’s own experiences in a technology such as an AI chatbot, allows for a ready template of comparison at the very least, and possibly an increase in ability to make sense of, and even predict, the agent’s behaviors. This may include being placed in a position to master a certain technology, whether by incentives to learn, or fear of poor outcomes should one not manage to learn.
@@
These pressures extend to AI experts, as well as laypersons. In both scholarly and commercial fields, AI experts face considerable pressure to demonstrate competence in their work. Citation metrics and scholarly publications remain the primary metric for tenure and promotion (Alperin et al., 2019), and the number of publications in the AI field has boomed as evidenced by overall (preprint and peer-reviewed) scholarly publications [^5] https://ourworldindata.org/grapher/annual-scholarly-publications-on-artificial-intelligence?time=2010..2021  (Maslej et al., 2023). The adoption of techniques underlying technologies with the AI label, i.e. machine learning, has spread to fields beyond Computer Science, e.g. Astronomy, Physics, Medicine and Psychology [^6] Retrieved 23/01/23 using the search string ”TITLE-ABS-KEY ( ( machine AND learning ) OR ( artificial AND intelligence ) OR ai ) AND PUBYEAR $>$ 2009 AND PUBYEAR $<$ 2024 ” from the SCOPUS database. Outside of academia, the number of jobs requiring AI expertise increases rapidly, with demand for ‘Machine Learning’ skills clusters having increased over 500% from 2010 to 2020 (Maslej et al., 2023). Thus, according to theory, the pressure to demonstrate AI-competence is fertile ground for anthropomorphism to occur.
@@ 4.3 Confirmation Bias
Hypotheses in present-day AI research are often implicit. Generally, these hypotheses are framed simply as a system being more accurate or efficient, compared to other systems. Where other fields, such as medicine or quantitative social sciences, would further articulate expectations in e.g. assigning specific conditions and considering effect sizes assigned to each competing hypothesis, in Computer Science and AI this is typically not done. This also may have to do with much of the published work being more of an engineering achievement, rather than a true hypothesis test seeking to explain and understand the world. However, in discussions on emerging qualities like AGI, this engineering positioning gets muddier, and more formal hypothesis testing would be justifiable: either one interprets outputs as in support of hints towards AGI (the alternative hypothesis), or as merely the result of an algorithm integrating qualities from the data it was trained on (the null hypothesis).
@@
5 Outlook
Reflecting on the previous two sections, we make the following concrete recommendations for future research:
  1. 1.
    (Acknowledgement of Human Bias) Researchers should be mindful of, and explicit about, risks of human bias and anthropomorphization in interpreting results, which both can be done as part of the results discussion, but also in a dedicated ‘limitations’ section.
  2. 2.
    (Stronger Testing) Researchers should refrain from drawing premature conclusions about AGI, unless these are based on strong hypothesis tests.
  3. 3.
    (Epistemologically Robust Standards) We call for more precise definitions of terms like ‘intelligence’ and ‘AGI’, and publicly accountable and collaborative iterations over how we will measure them, with explicit room for independent reviewing and external auditing by the broader community.
Moreover, we believe that structural and cultural changes are in order to reduce current incentives to chase Big Statement Outcomes in AI research and industry. Our broadest and perhaps most ambitious goal is for our research community to move away from authorship and instead embrace contributorship. This argument has been raised long before in other research communities (Smith, 1997) and more recently within our own (Liem & Demetriou, 2023). Specifically, Liem & Demetriou (2023) argue that societally impactful scientific insights should be treated as open-source software artifacts. The open-source community sets a positive example of how scientific artifacts should be published in many different ways. Not only does it adequately reward small contributions but it also naturally considers negative results (bugs) as part of the scientific process. Similarly, code reviews are considered so integral to the process that they typically end up as accredited contributions to projects. Open review platforms like OpenReview are a step in the right direction, but still fall short of what we know is technically feasible. Finally, software testing is, of course, not only essential but often obligatory before contributions are accepted and merged. As we have pointed out repeatedly in this work, any claims about AGI demand proper strong hypothesis tests. It is important to remember that AGI remains the alternative hypothesis and that the burden of proof therefore lies with those making strong claims.
@@ 6 Conclusion
As discussed above, AI research and development outcomes can easily be over-interpreted, both from a data perspective and because of human biases and interests. Academic researchers are not free from such biases. Thus, we call for the community to create explicit room for organized skepticism.
@@
For research that seeks to explain a phenomenon, clear hypothesis articulation and strong null hypothesis formulation are needed. If claims of human-like or superhuman intelligence are made, these should be subject to severe tests (Claesen et al., 2022) that go beyond the display of surprise. Apart from focusing on getting novel improvements upon state-of-the-art published, organizing red-teaming activities as a community may help in incentivizing and normalizing constructive adversarial questioning. As the quest for AGI is so deeply rooted in human-like recognition, adding our voice to emerging calls to be vigilant in communication (Shanahan, 2024), we put in an explicit word of warning about the use of terminology. Many terms used in current AGI research (e.g. emergence, intelligence, learning, ‘better than human’ performance) have a common understanding in specialized research communities, but have bigger, anthropomorphic connotations in laypersons. In fictional media, depictions of highly intelligent AI have for long been going around. In a study of films featuring robots, defined as ”…an artificial entity that can sense and act as a result of (real-world or fictional) technology…”, in the 134 most highly rated science-fiction movies on IMDB, 74 out of the 108 AI-robots studied had a humanoid shape, and 68 out of those had sufficient intelligence to interact at an almost human-level (Saffari et al., 2021). The authors identify human-like communication and the ability to learn as essential abilities in the depiction of AI agents in movies. They further show a common plot: humans perceive the AI agents as inferior, despite their possession of self-awareness and the desire to survive, which fuels the central conflict of the film, wherein humanity is threatened by AI superior in both intellect and physical abilities. It is often noted that experts and fictional content creators interact, informing and inspiring each other (Saffari et al., 2021; Neri & Cozman, 2020).
@@
This image also permeates present-day non-fictional writings on AI, which often use anthropomorphized language (e.g. “ever more powerful digital minds” in the ‘Pause Giant AI Experiments’ open letter (Future of Life Institute, 2023)). In the news, we witness examples of humans falling in love with their AI chatbots (Morrone, 2023; Steinberg, 2023). The same news outlets discuss the human-like responses of Microsoft’s Bing search engine, which had at that point recently been adopting GPT-4777A large multimodal language model from OpenAI https://openai.com/research/gpt-4. The article (Cost, 2023), states “As if Bing wasn’t becoming human enough” and goes on to claim it told them it loves them. Here, AI experts and influencers also have considerable influence on how the narrative unfolds on social media: according to Neri & Cozman (2020), actual AI-related harms did not trigger viral amplification (e.g. the death of an individual dying while a Tesla car was in autopilot, or the financial bankruptcy of a firm using AI technology to execute stock trades). Rather, potential risks expressed by someone perceived as having expertise and authority were amplified, such as statements made by Stephen Hawking during an interview in 2014.
@@
Acknowledgements
Some of the members of TU Delft were partially funded by ICAI AI for Fintech Research, an ING—TU Delft collaboration.
CRediT Author Statement
Following the CRediT Contributor Roles Taxonomy, we list author contributions in descending order of degree of contribution: Conceptualization PA, AD, AB, CL; Data curation PA; Formal analysis PA; Funding acquisition CL; Investigation PA; Literature review AD, AB, CL; Methodology PA, CL; Project administration PA, AD, AB; Software PA; Supervision CL; Visualization PA; Writing - original draft PA, AD, AB; Writing - review and editing PA, AD, AB, CL.
@@ A.2.1 Data
To estimate economic growth, we will rely on a quarterly series of the real gross domestic product (GDP) provided by the Federal Reserve Bank of St. Louis. The data arrives in terms of levels of real GDP. In order to estimate growth, we transform the data using log differences. Since our yield curve data is daily, we aggregate it to the quarterly frequency by taking averages of daily yields for each maturity. We also standardize yields since deep learning models tend to perform better with standardized data (Michal S Gal, 2019). Since COVID-19 was a substantial structural break in the time series, we also filter out all observations after 2018.
@@ A.2.2 Model
The encoder consists of a single fully connected hidden layer with 32 neurons and a hyperbolic tangent activation function. The bottleneck layer connecting the encoder to the decoder, is a fully connected layer with 6 neurons. The decoder consists of two fully connected layers, each with a hyperbolic tangent activation function: the first layer consists of 32 neurons and the second layer will have the same dimension as the input data. The output layer consists of a single neuron for our output variable, $g_{t}$. We train the model over 1,000 epochs to minimize mean squared error loss using the Adam optimizer (Kingma & Ba, 2017).
@@ A.2.3 Linear Probe
While the loss function applies most direct pressure on layers near the final output layer, any information useful for the downstream task first needs to pass through the bottleneck layer (Alain & Bengio, 2016). On a per-neuron basis, the pressure to distill useful representation is therefore likely maximized there. Consequently, the bottleneck layer activations seem like a natural place to start looking for compact, meaningful representations of distilled information. We compute and extract these activations $A_{t}$ for all time periods $t=1,...,T$. Next, we use a linear probe to regress the observed yield curve factors on the latent embeddings. Let $Y_{t}$ denote the vector containing the two factors of interest in time $t$: $y_{t,l}$ and $y_{t,s}$ for the level and spread, respectively. Formally, we are interested in the following regression model: $p_{w}(Y_{t}|A_{t})$ where $w$ denotes the regression parameters. We use Ridge regression with $\lambda$ set to $0.1$. Using the estimated regression parameters $\hat{w}$, we then predict the yield curve factors: $\hat{Y}_{t}=\hat{w}^{\prime}A_{t}$.